Presentation on theme: "Lau Gatignon, EATM, 7 October 2014. The T9 and T10 lines have been running smoothly from mid July and the T11 beam started up very smoothly for CLOUD."— Presentation transcript:
Lau Gatignon, EATM, 7 October 2014
The T9 and T10 lines have been running smoothly from mid July and the T11 beam started up very smoothly for CLOUD as scheduled. The DSO test for CHARM and IRRAD had to be postponed by week until October 3 rd due to a (voluntarily cut cable in the access system (see pictures). Two rollers supporting the mobile shielding have failed, due to a non- conformity in the production process (the specified thermal treatment was not done). New rollers have been ordered, but they will only arrive in a few weeks from now, In the meantime tests with one low-intensity spill per super-cycle can start on Thursday 9 th of October, in agreement with RP. The ventilation system of CHARM and IRRAD is expected to become operational by the end of October. Once this is ok and the new rollers have been installed, operation at nominal intensity can start. Thanks for all the efforts that made this installation possible.
The cut cable
T9:CLIC-Pix continues till Wednesday 8 th of October. Then change to CALICE (W.Klempt), who have prepared their calorimeter downstream of T9 at the downstream end of the hall. On Wednesday 22 nd of October change to FCAL (L.Linssen), a small e.m. calorimeter (< 200 kg). They may use the EUDET telescope during the first part of the run. On Wednesday 29 th of October change to DAMPE. They require the installation of a MDX magnet inside the T9 zone (in the traditional location of MNP17, this has been done in the past). Otherwise mainly Silicon detectors and a calorimeter. They also request a DESY table and Iron blocks to serve as muon filter. A schematic layout is shown in the next slide.
T10:Today change to ALICE FIT T0, a well known user with a small detector with a quartz radiator + MCP-PMT. On Tuesday 14 th of October change to COMPASS ECAL. A test of a shshlik calorimeter, 200 kg. They request the Cerenkov counter. On Friday 24 th of October change to ALICE ITS, a small detector on an existing support. Remove the threshold Cerenkov. Om Friday October 31 st change to a combination of ALICE tests, all on the existing ALICE infrastructure. T11: CLOUD continues, no changes
The double wall is installed and equipped in TCC8/ECN3 and will be tested during the scrubbing run. The air lock is still missing in GL300. This is ok as only low intensities will be used until then. The beam line has been completed and the setting up is starting. Also the detector is complete except MUV1 and GTK1. The latter will be installed very soon. It is important that P0 survey and DUMP CONTROL are operational. Many thanks for all the work done by all of you!
Start with muon beam immediately after the EATM meeting. The CEDAR leaks have been ‘fixed’ and they are now tight enough. The BE absorbers needed quite some effort to make them operational again (oxydation). This will require some more thorough maintenance in the next shutdown.